OpenStudy (anonymous):

What fractional part of a degree is 45'? __ / __ °

OpenStudy (anonymous):

Root 2 over 2

OpenStudy (anonymous):

how do you get the answer?

OpenStudy (anonymous):

well we draw a diangonal through a square, with sides 1, then using the pytagorean throem on that we get the hypotenuse which will be root 2

OpenStudy (anonymous):

According to me, the answer is different 1 degree = 60' so 45'/60' gives 3/4 after simplification

OpenStudy (anonymous):

What answer do u have??

OpenStudy (anonymous):

or the answer could also show up as 1over root 2, after rationalizing

OpenStudy (anonymous):

\[\frac{\sqrt{2}}{2}\]is the sine and cosine of a 45 degree angle. it is not a 'fractional part of 45" whatever that means

OpenStudy (anonymous):

i think this means "what part of a degree is 45 minutes" in which case, since there are 60 minutes in one degree the portion is \[\frac{45}{60}\]

OpenStudy (anonymous):

or if you prefer \[\frac{3}{4}\] meaning 45 minutes is 3/4th of one degree

OpenStudy (anonymous):

it was \[45'\] yes

OpenStudy (anonymous):

then Harkirat had it